OpenStudy (anonymous):

karen has 4h storybooks . Karens storybook collection is 2/3 of what Mary's collection is

OpenStudy (superflytyguy):

4h as in 400?

OpenStudy (superflytyguy):

If so... 2/3 = 0.666 (6 never ends) Multiply 2/3 x 400 2/3 x 400 = 266